OpenStudy (anonymous):

A triangle has side lengths of 33, 56, and 64. It is a right triangle? Explain yes; 33^2 + 56^2 = 65^2 yes; 33^2 + 56^2 ≠ 65^2 no; 33^2 + 56^2 = 65^2 no; 33^2 + 56^2 ≠ 65^2

OpenStudy (anonymous):

@ganeshie8 can u help

OpenStudy (anonymous):



OpenStudy (anonymous):

condition to be a right triangle h²= a²+b²

OpenStudy (anonymous):

then... no; 33^2 + 56^2 = 65^2

OpenStudy (anonymous):

@Compassionate

OpenStudy (radar):

Be careful, Did you actually square those sides and sum them to see if they equaled the hypotenuse squared?

OpenStudy (radar):

\[33^2 + 56^2 =1089 + 3136 = 4225\]

OpenStudy (radar):

What does 65^2 equal ??

OpenStudy (anonymous):

4225

OpenStudy (radar):

If 65^2 = 4225, then yes

OpenStudy (anonymous):

so why would it be yes

OpenStudy (radar):

because the sum of the squared legs equal the square of the hypotenuse..... Pythagorean says so for the right triangle. The longest side is the hypotenuse.

OpenStudy (anonymous):

so u saying that the answer is A

OpenStudy (radar):

I am saying you should say A is the answer.

OpenStudy (anonymous):

im not understanding you is a the answer or not

OpenStudy (radar):

You figure it out. does 1089 + 3136 = 4225 Did you not read my posts above??

OpenStudy (radar):

What is answer A. saying??